OpenStudy (shesolitt):

I WILL MEDAL AND FAN What is the equation, in slope-intercept form, of the line that passes through (0, 3) and has a slope of −6? y = 3x + 6 y = 3x − 6 y = −6x − 3 y = −6x + 3

pooja195 (pooja195):

Hi there! What do you think the answer is and why?

OpenStudy (shesolitt):

@dumptybaby @jackthegreatest @jhonyy9 @kidrah69 @Rz172

OpenStudy (shesolitt):

i think its c

pooja195 (pooja195):

Well the 3 in the coordinate is not negative

OpenStudy (jackthegreatest):

the point (0,3) means that when x = 0, y = 3

OpenStudy (jackthegreatest):

if you used C, y = −6x - 3 3 = - 6(0) - 3 3 does not equal -3

OpenStudy (shesolitt):

oh okay. then what? lol

OpenStudy (jackthegreatest):

try the same thing for D

OpenStudy (jackthegreatest):

show me your work

OpenStudy (shesolitt):

okay

OpenStudy (shesolitt):

y = −6x + 3 3 = - 6(0) + 3

OpenStudy (jackthegreatest):

yep, and do the two sides equal?

OpenStudy (shesolitt):

yes

OpenStudy (shesolitt):

so the answers D

OpenStudy (jackthegreatest):

correct : )

OpenStudy (shesolitt):

thank you (:

OpenStudy (jackthegreatest):

np